Reports 1-1 of 1 Clear search Modify search
AdV-TCS (TCS control system)
mantovani, nardecchia - 17:19 Thursday 19 May 2022 (55908) Print this report
Warm CH and DAS tuning using cavity matching technique

This morning we have worked on the TCS tuning for the Warm.

Starting configuration: DAS and CH off (initial data points in Figure 1) mode matching value 0.0148.

@ 8:10UTC the WI CH power has been set to 93mW to mimic the YAG beam (~0.572W on the pick-off powermeter), which corresponds to the beginning of the first yellow circles in Figure 1.

@ 9:46UTC the WI DAS has been turned on (0.4W IN 1.1W OUT, values expected from simulations). A bit before than the light cyan circles, see Figure 1.

The Figure 1 shows the matching evolution starting from 7:38:30 UTC. The 00 mode has been measured @ 7:54:40UTC and @ 14:06:15UTC giving respectively : 23681.7038,  23637.409 (values to use for the 02Mode normalization).

It is visible that after the DAS compensation the matching is back to the initial value, which means that the 0 lens is achieved.

It is visible that the matching is well in agreement with the circulating power in the cavity, see Figures 2 and 3. (the last data points are not compared because the agreement was not so optima, which is most probably due to etalon effect).

Side note:

- the north cavity was badly misaligning during the scan (with drift control off). It has been improved by engaging the Drift control on the cavity mirror. Same situation for some data set of West cavity.

The command to act on the drift control (no PR translation and BS tilt) is drift_control_TMs(1) to engage and drift_control_TMs(0) to disengage.

Images attached to this report
Search Help